OpenStudy (anonymous):

can you please check if i got this right also help me with one question as soon as possible pleaasee

OpenStudy (anonymous):

4. Use the following arithmetic sequence and the formula to answer the questions below. 123, 116, 109, 102, 95, ... Part I: Find the value of each of the following: a1 = _123______, a2 = __116_____, and d = ___ a2-a1 = 116-123 = -7 ____. (3 points) Part II: Find the explicit formula. Show your work. an = a1 + d(n-1) = 123 - 7(n-1) = 123 -7n + 7 = 130 - 7n a100 = 130 - 7*100 = 130 - 700 = -570 Part III: Use the explicit formula you found in Part II to find the value of the 100th term in the sequence, a100. Show your work.

OpenStudy (anonymous):

i dont know how to do part 3

OpenStudy (anonymous):

also im not sure if i got part 2 right either

OpenStudy (popitree):

u got the right formula, nth no = 130 - 7n. so if u want to find 100th no a100 = 130 - 7.100

OpenStudy (anonymous):

so i got evreything right and for part 3 is 130-7.100 right

OpenStudy (popitree):

yes

OpenStudy (anonymous):

thank you sooo much